The concentration of a reactant X decreases from 0.1 M to 0.025 M in 40 minutes. If the reaction follows I order kinetics, the rate of the reaction when the concentration of X is 0.01 M will be

(A) 1.73 ´ 10-4 M min-l
(B) 2.47 ´ 10-4 M min-l
(C) 3.47 ´ 10-5 M min-l
(D) 1.73 ´ 10-5 M min-l

Grade:12

1 Answers

Nirgun
4 years ago
k=2.303/t log(a/a-x)
a=0.1M
a-x=0.025M
t=40min
 
 
k=2.303/40 log (0.1/0.025)
 =0.0347^-1 min^-1
 
Rate as it is 1st order
Rate  dx/dt=R[a]^1
                 = 0.0347*0.01
                 =3.4*10^-4 Mmin^-1